OpenStudy (anonymous):

Find tan of angle B. Write your answer as a simplified fraction. Use the triangle below to answer this question.

OpenStudy (anonymous):



OpenStudy (anonymous):

I found Angle A=53 degrees Ange B=37 Degrees and side b=15 PLEASE HELP

OpenStudy (raden):

first, can u find the length of b by using pythagorean theorem ?

OpenStudy (raden):

nevermind.. u have got it... :) now, tanB = opp/adj = 15/20 = 3/4